A diver of weight 713 N standing at the end of a diving board with a length L = 4.1 m and negligible mass. The board is fixed to two pedestals (supports) separated by distance d = 1.5 m.

(a) Of the forces acting on the board, what is the magnitude of the force from the left pedestal? N

(b) Of the forces acting on the board, what is the direction of the force from the left pedestal? up or down?

(c) Of the forces acting on the board, what is the magnitude of the force from the right pedestal? N

(d) Of the forces acting on the board, what is the direction of the force from the right pedestal? up or down?

(e) Which pedestal is being stretched?

- The left pedestal is being stretched.

- The right pedestal is being stretched.

(f) Which pedestal is being compressed?

- The left pedestal is being compressed.

- The right pedestal is being compressed.
Answer #1

a)The force acting on the board is, F/+F, -W = 0 FD+W (L-D)=0 -(L-D). D -(4.1-1.5) (713) 1.5 =-1235.866667 =-1236N |FI=1236N